OpenStudy (anonymous):

Simplify the following using scientific notation:  (8.76 x 109)(6.52 x 10-3)/13.27 x 105

OpenStudy (ybarrap):

Is it $$ \cfrac{(8.76 \times 109)(6.52 \times 10^{-3})}{13.27 \times 105} $$ Or $$ \cfrac{(8.76 \times 109)(6.52 \times 10^{-3})}{13.27 }\times 105 $$ ?